Circular No. 6450

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

4U 0115+634 M. Scott and M. H. Finger, Universities Space Research Association; R.

B. Wilson, Marshall Space Flight Center, NASA; and T. A. Prince and B. Vaughan, California Institute of Technology, report for the Compton GRO Observatory BATSE pulsar team: “Pulsed hard x-ray flux is currently being detected from the transient Be/x-ray pulsar 4U 0115+634. The source was detected on Aug. 10, when the rms pulsed flux (20–50 keV) was ∼ 25 mCrab. This is the first detection by BATSE of this source since a pair of outbursts that occurred during the interval 1995 Nov. 18–1996 Jan. 9 (IAUC 6266). On Aug. 12, the rms pulsed flux (20–50 keV) was ∼ 30 mCrab. The pulse frequencies observed from Aug. 10 to 12 are consistent with a neutronstar spin rate of 0.27666712 ± 0.00000022 Hz and no detectable spin-up, using the binary orbital ephemeris of Rappaport et al. (1978, Ap.J. 211, L1) with the epoch of periastron passage updated to 1996 Aug. 5.5 UT.”

NOVA SAGITTARII 1996 L. V. Morrison and R. W. Argyle, Royal Greenwich Observatory, re-

port an accurate optical position for N Sgr 1996, from nine observations obtained with the Carlsberg Automatic Meridian Circle on La Palma and operated by C. Fabricius and L. V. Morrison: α = 18^{h}23^{m}42^{s}.501, ± 0^{s}.005, δ = −18^{o}07′14′′.80 ± 0′′.07 (equinox J2000.0, FK5 reference frame, epoch 1996.61). The following V magnitudes have also been measured (±0.08 mag): Aug. 5.94 UT, 11.01; 6.94, 11.20; 7.94, 11.54; 8.93, 11.65; 9.93, 11.96; 10.93, 12.29; 11.92, 12.07; 12.92, 12.19; 13.92, 12.15.

E. Martin, Instituto de Astrofísica de Canarias, reports the following near-infrared magnitudes obtained with the 1.5-m Carlos Sanchez telescope (+ CVF photometer) at Teide Observatory on Aug. 9.965 (heliocentric) UT: J = 8.93 ± 0.04, H = 8.73 ± 0.03, K = 8.35 ± 0.03.

PERSEID METEORS 1996 A. Kereszturi, K. Sárneczky, and I. Tepliczky, Hungarian Astronomical

Association; and P. Jenniskens, NASA Ames Research Center, report that observations from Europe suggest that the peak in 1996 Perseid activity occurred around Aug. 12.06 UT, at up to two or three times the normal rates. C. M. Gaskell, University of Nebraska, writes that the mean Perseid magnitude on Aug. 12 was almost two mag brighter than historic means.

1996 August 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green